OpenStudy (anonymous):

If y varies jointly as x and z and y=80 when x=5 and z=8, find y when x=16 and z=2

OpenStudy (diyadiya):

y varies jointly so y=k * xz

OpenStudy (diyadiya):

80=k *5*8

OpenStudy (anonymous):

and jus solve for k?

OpenStudy (diyadiya):

Yeah find out the value of k first

OpenStudy (anonymous):

okie dokie!! thanks diya!!!

OpenStudy (diyadiya):

then you've to substitute it in y= k *xz for the new value of x & z and find out y

OpenStudy (anonymous):

yea, okay, gotcha!! thank you thank you!!!!

OpenStudy (diyadiya):

hehe :) Anytime !!